OpenStudy (anonymous):

MEDAL AND FAN FOR ANSWER Paula has a large ball of yarn that she is going to use to knit a scarf for the winter. Every square inch on the scarf requires a certain number of yards of yarn from the ball. A linear model of this situation contains the values (18, 240.72) and (30, 97.2), where x represents the number of square inches knitted on the scarf, and y represents the number of yards remaining on the ball of yarn. How many yards of yarn did Paula start with?

OpenStudy (anonymous):

translate the question as "find the equation of the line between the points \[(18,240.72)\] and \[(30,97.2)\] when it is in the form \(y=mx+b\) the answer will be \(b\) for what you started with

OpenStudy (anonymous):

i can walk you through it if you like do you know how to find the equation of a line given two points?

OpenStudy (anonymous):

a.504 b.11.96 c.337.92 d.456 @satellite73

OpenStudy (anonymous):

go with D

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

equation of the line needs the slope slope is \[\frac{97.2-240.72}{30-10}=-11.96\] (used a calculator)

OpenStudy (anonymous):

then the point slope formula gives \[y-97.2=11.96(x-30)\]

OpenStudy (anonymous):

456 yards

OpenStudy (anonymous):

typo there \[y-97.2=- 11.96(x-30)\]

OpenStudy (anonymous):

solve this equation for \(y\) and get \[y=-11.96x+456\]

OpenStudy (anonymous):

what you started with is what you get when \(x=0\) in other words the \(y\) intercept which is \(456\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you. it was 456

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ur in which grade jessica

OpenStudy (anonymous):

lol said 5 minutes ago did you need all that explanation?